# R3R: Decentralized Multi-Agent Collision Avoidance with Infinite-Horizon Safety

### v1

@@ Abstract
Existing decentralized methods for multi-agent motion planning lack formal, infinite-horizon safety guarantees, especially for communication-constrained systems. We present R3R, to our knowledge the first decentralized and asynchronous framework for multi-agent motion planning under distance-based communication constraints with infinite-horizon safety guarantees for systems of nonlinear agents. R3R’s novelty lies in combining our gatekeeper safety framework with a geometric constraint called R-Boundedness, which together establish a formal link between an agent’s communication radius and its ability to plan safely. We constrain trajectories to within a fixed planning radius that is a function of the agent’s communication radius, which enables trajectories to be shown provably safe for all time, using only local information. Our algorithm is fully asynchronous, and ensures the forward invariance of these guarantees even in time-varying networks where agents asynchronously join, leave, and replan. We validate our approach in simulations of up to 128 Dubins vehicles, demonstrating 100% safety in dense, obstacle rich scenarios. Our results demonstrate that R3R’s performance scales with agent density rather than problem size, providing a practical solution for scalable and provably safe multi-agent systems [^1] Code available at github.com/MarshallVielmetti/r3r_decentralized_gatekeeper.
@@ I Introduction
From last mile delivery [1] to warehouse robotics [2] to autonomous drone swarms [3], deploying multi-agent systems safely in the real world is bottlenecked by a lack of formal safety guarantees. This challenge, known as multi-agent motion planning (MAMP), forces a trade-off between guarantees, coordination, and scalability. Centralized planners can provide strong safety guarantees but fail to scale and are prone to single-node failure. Conversely, decentralized methods scale well to large teams but have thus far been unable to provide formal, infinite-horizon safety guarantees under distance-based communication constraints for non-trivial (i.e. beyond single/double integrator) systems.
Multi-agent path finding (MAPF), a discrete state precursor to MAMP, is typically formulated as a graph search problem. Classical MAPF methods, such as Conflict-Based Search and its variants [4, 5, 6, 7], provide strong optimality or bounded suboptimality guarantees, but these approaches are centralized and limited to discrete domains, thus may not produce dynamically feasible solutions. Some MAMP methods have extended these concepts to continuous space, and trajectory optimization and sampling-based planners have been proposed, but these centralized approaches struggle with scalability and real-time performance [8].
@@
Critically, no existing decentralized, asynchronous approach guarantees infinite-horizon safety under distance-based communication constraints. This gap motivates our proposed framework, R3R. To the best of our knowledge, R3R represents the first decentralized and asynchronous framework to provide infinite-horizon safety guarantees for multi-agent motion planning under distance-based communication constraints, and generalize to broad classes of systems under certain assumptions. This is achieved by integrating a novel geometric constraint, which we term R-Boundedness with gatekeeper [18], our safety framework developed for single agents, to provide recursive feasibility guarantees in multi-agent systems. Our contributions are as follows:
@@
    R3R, a decentralized and asynchronous framework for multi-agent motion planning under communication constraints that provides infinite-horizon safety guarantees.
@@
    The key enabler of R3R is R-Boundedness, a novel geometric constraint on trajectory generation, which we show links an agent’s ability to communicate with its ability to plan safely.
@@
    We demonstrate the efficacy of R3R in simulations of up to 128 agents, showcasing safe, and scalable multi-agent coordination.
The paper is organized as follows: Section II introduces preliminaries and the problem statement, and section III presents the proposed approach. The methodology is evaluated in simulation in section IV, while conclusions and thoughts for future work are given in section V.
@@ II Preliminaries and Problem Statement
Notation: $\mathbb{R},\mathbb{R}_{\geq 0},\mathbb{R}_{>0}$ are the sets of reals, non-negative reals, and positive reals. Let $\mathcal{B}(x,R)=\{y:\left\|x-y\right\|\leq R\}$ be the closed ball of radius $R\geq 0$ centered at $x\in\mathbb{R}^{n}$. $\mathcal{A}(t)=\{1\dots N\}$ denotes the set of all agents at time $t$, where $N$ is the total number of agents in the network at that time. Let $\mathcal{A}_{-i}(t)=\mathcal{A}(t)\setminus\{i\}$ for $i\in\mathcal{A}(t)$. Let $\delta\in\mathbb{R}_{>0}$ denote the inter-agent collision radius.
@@ II-A Preliminaries
where $x\in\mathcal{X}\subseteq\mathbb{R}^{n}$ is the state, $u\in\mathcal{U}\subseteq\mathbb{R}^{m}$ is the control input. The function $f:\mathbb{R}_{\geq 0}\times\mathcal{X}\times\mathcal{U}\to\mathbb{R}^{n}$ is piecewise continuous and locally Lipschitz in $x$ and $u$.
@@
    The agent’s state $x_{i}$ evolves through a system satisfying (1), and includes a position component $p\in\mathbb{R}^{d}$.
@@
    $\delta\in\mathbb{R}_{>0}$ Denotes the inter-agent avoidance distance. No other agent may come within a distance $\delta$ of agent $i$.
@@
    $R^{\text{comm}}\in\mathbb{R}_{>0}$ represents the communication radius.
  4. 4.
    A planning radius $R^{\text{plan}}$ which constrains an agent’s future trajectory.
This relationship between $R^{\text{comm}}$ and $R^{\text{plan}}$, which we show must be $R^{\text{comm}}=3R^{\text{plan}}+\delta$, (R3R) is the key insight of this work, and will be explored in detail.
@@
All agents have a common $\delta$ and $R^{\text{comm}}$, and thus a common $R^{\text{plan}}$.
@@
A trajectory [^2] The time-interval $T$ can be a finite horizon, e.g., $T=[t_{0},t_{1}]$ or an infinite horizon, e.g., $T=[t_{0},\infty)$. is the tuple $\mathcal{T}=(T,x,u)$, where $T\subset\mathbb{R}_{\geq 0}$ is a time-interval, $x:T\to\mathcal{X}$ is the state-trajectory, and $u:T\to\mathcal{U}$ is a control-trajectory, satisfying
$\dot{x}(t)=f(t,x(t),u(t)),\quad\forall{}t\in T.$ (2)
Definition 3.
Let $\mathcal{G}(t)=(\mathcal{A}(t),E(t))$ be an undirected, time-varying graph, with nodes representing agents, and edges representing communication links between them. The edge set $E(t)$, defined for some communication radius $R^{\text{comm}}>0$, is given by:
$E(t):=\{(i,j)\in\mathcal{A}(t):\left\|p_{i}(t)-p_{j}(t)\right\|\leq R^{\text{comm}}\}$ (3)
where $p_{i}(t)\in\mathbb{R}^{d}$ is the position of agent $i\in\mathcal{A}(t)$ at time $t$.
@@
We assume the communication is instantaneous and range-limited, but not bandwidth limited. That is, if two agents $i,j$ are neighbors at some time $t$, they can communicate an arbitrary amount of information instantaneously.
We define the set of neighbors of agent $i$ at time $t\in\mathbb{R}_{\geq 0}$ as $\mathcal{N}_{i}(t)$, where
$\mathcal{N}_{i}(t)=\{j\in\mathcal{A}(t):(i,j)\in E(t)\}.$ (4)
Definition 4 ($\operatorname{dist}$).
Let $\mathcal{S}_{1},\mathcal{S}_{2}\subset\mathbb{R}^{d}$, $x\in\mathbb{R}^{d}$, define
$\operatorname{dist}(\mathcal{S}_{1},x)$ $=\inf_{y\in\mathcal{S}_{1}}\left\|x-y\right\|_{2},$ (5)
$\operatorname{dist}(\mathcal{S}_{1},\mathcal{S}_{2})$ $=\inf_{x\in\mathcal{S}_{1},y\in\mathcal{S}_{2}}\left\|x-y\right\|_{2}.$ (6)
II-B gatekeeper Preliminaries
R3R builds upon our prior work for single-agent safety gatekeeper [18], which ensures that agents always track a committed trajectory that is known to be safe for all future time. This is done by properly appending the goal-oriented nominal trajectory with a safe infinite-horizon backup trajectory. This becomes a candidate trajectory, which must be validated as safe prior to being committed. The steps are discussed later, and the principle is shown in fig. 1. Our key extension is to impose an additional geometric constraint on candidate trajectories, as illustrated in fig.˜1, which enables decentralized verification.
@@
Let $\mathcal{S}\subset\mathcal{X}$ be the safe set, to be defined in ˜1.
@@
A set $\mathcal{C}\subset\mathcal{X}$ is a backup set if $\mathcal{C}\subset\mathcal{S}$, and there exists a controller $\pi^{B}:\mathbb{R}\times\mathcal{X}\to\mathcal{U}$ such that $\forall t_{i}\in\mathbb{R}_{\geq 0}$, the closed loop system $\dot{x}=f(t,x,\pi^{B}(t,x))$ satisfies
$x(t_{i})\in\mathcal{C}\implies x(t)\in\mathcal{C},\forall t\geq t_{i}.$ (7)
Definition 6 (Nominal Trajectory).
Given state $x(t_{k})\in\mathcal{X}$ at some time $t_{k}\in\mathbb{R}_{\geq 0}$, a nominal trajectory, typically generated by a planner, is denoted
$(T_{k}^{\rm nom},x_{k}^{\rm nom},u_{k}^{\rm nom})$ (8)
and defined over the horizon $T_{k}^{\rm nom}=[t_{k},t_{k}+T_{H}]$, where $T_{H}\in\mathbb{R}_{>0}$ is the planning horizon.
A nominal trajectory is not guaranteed to satisfy constraints, nor guarantee safety beyond $T_{H}$. In the context of communication constrained planning, this is because nominal trajectories are constructed without full knowledge of the system state, and thus may collide with agents outside the planning agents communication range.
@@
Let $T_{B}\in\mathbb{R}_{>0}$. For any $t_{s}\in\mathbb{R}_{>0}$, and $x_{s}\in\mathcal{X}$, a trajectory $(T^{\rm bak},x^{\rm bak},u^{\rm bak})$, defined on $T^{\rm bak}=[t_{s},\infty)$ and terminating in a backup set $\mathcal{C}$ is a backup trajectory if
$x^{\rm bak}(t_{s}+T_{B})\in\mathcal{C},$ (9)
and for all $t\geq t_{s}+T_{B}$,
$u^{\rm bak}(t)=\pi^{B}(t,x^{\rm bak}(t)).$ (10)
A backup trajectory takes a system from a state $x\in\mathcal{X}$ at time $t_{s}$ such that at time $t_{s}+T_{B}$, the system lies in a backup set. The system then executes the backup controller for all future time, thus remaining in the backup set. If the finite horizon trajectory defined over $t_{s},t_{s}+T_{B}]$ safely takes the agent into the backup set, then by the properties of the backup set we guarantee the agent will be safe over $t_{s},\infty)$, transforming a finite horizon guarantee into an infinite horizon one.
@@
At time $t_{k}\in\mathbb{R}_{\geq 0}$, let agent $i\in\mathcal{A}(t)$ be at state $x_{i}(t_{k})\in\mathcal{X}$. Let the agent’s nominal trajectory be $\mathcal{T}_{k,i}^{\rm nom}=([t_{k},t_{k}+T_{H}],x_{k,i}^{\rm nom},u_{k,i}^{\rm nom})$. A candidate trajectory with switch time $t_{s}\in[t_{k},t_{k}+T_{H}]$, denoted $\mathcal{T}_{k,i}^{{\rm can},t_{s}}=([t_{k},\infty),x_{k,i}^{\rm can},u_{k,i}^{\rm can})$ is defined by:
$\mathcal{T}_{k,i}^{\rm can}=\begin{cases}(x_{k,i}^{\rm nom}(\tau),u_{k,i}^{\rm nom}(\tau))&\text{if }\tau\in[t_{k},t_{s})\\ (x_{k,i}^{\rm bak}(\tau),u_{k,i}^{\rm bak}(\tau))&\text{if }\tau\geq t_{s}\end{cases}$ (11)
where $(x_{k,i}^{\rm bak},u_{k,i}^{\rm bak})$ is a backup trajectory from $t_{s},x_{k}^{\rm nom}(t_{s})$, as in definition˜7.
II-C Problem Statement
Problem 1.
Subject to assumptions˜1 and 2, let $\mathcal{A}(t)=\{1,2,\dots,N\}$ denote a time-varying set of agents satisfying definition˜1. These agents communicate over a connectivity graph $\mathcal{G}(t)$ defined in definition˜3. Let $\mathcal{X}_{\text{obs}}\subset\mathcal{X}$ be the subset of the domain occupied by static obstacles, and $\mathcal{S}\subset\mathcal{X}\setminus\mathcal{X}_{\text{obs}}$ be the safe set. Assume this set is known to all agents. Let the goal state of each agent be denoted as $g_{i}\in\mathcal{S}$. Agents attempt to navigate to their goal state, such that
$\lim_{t\to\infty}x_{i}(t)=g_{i},\quad\forall i\in\mathcal{A},$ (12)
while satisfying the static and inter-agent collision avoidance constraints
$x_{i}(t)\in\mathcal{S},$ $\quad\forall t\in\mathbb{R}_{\geq 0},$ (13)
[…]
To solve ˜1, we propose R3R, a decentralized framework for multi-agent motion planning under communication constraints. R3R combines the gatekeeper framework with the concept of $R$-bounded trajectories, which we introduce in the next section. We then present a geometric argument that allows an agent to certify its trajectory as safe for all future time using only local information, and leverage gatekeeper to ensure forward invariance.
Figure 2: Demonstration of R3R over time for two Dubins agents. The timeline indicates when each agent (Red: Agent 1, Blue: Agent 2) plans or replans. Committed trajectories (solid lines) are certified to be safe for all time, and remain within a planning radius $R^{\text{plan}}$ (filled circles), around where they are constructed. Agents first plan goal-oriented nominal trajectories (dashed lines). A portion of this nominal is combined with a backup trajectory, to form a candidate. If this candidate is deemed safe, it becomes a committed trajectory. At times $t_{0,1}\neq t_{0,2}$ agents $1$ and 22 wish to join the network. Each agent independently constructs a valid committed trajectory, which is then simultaneously committed and broadcast as the agent enters the network. (a.) The initial committed trajectories of both agents remain within the planning radius. The agents could not communicate at the time of planning, so we see their nominal trajectories, which may be unsafe, intersect. (b.) Both agents have continued to track their initial committed trajectories, and are now within communication range. Both agents are near their backup sets, and thus should soon replan to continue making progress. (c.) Agent 11’s logic dictated performing a replan first, so at time $t_{1_{1}}$, it constructs a new nominal trajectory to its goal, which attempts to avoid agent $2$. It then constructs a new committed trajectory, which it certifies as safe with respect to the committed trajectory of agent $2$. At some time $t_{1_{2}}>t_{1_{1}}$, agent 22 constructs a new committed trajectory which must avoid the recently committed trajectory of agent $1$. This reflects the asynchronous nature of the algorithm, where agents replan independently and assume a low priority with respect to all neighbors. (d-f.) The process continues throughout the encounter. We see that agent 22 replans again at time $t_{2_{2}}$, but as agent $1$ has not yet reached its backup set, its internal logic does not necessitate a replan.
III Proposed Approach and Solution
We first present $R$-Boundedness, the theoretical component which allows us to ensure inter-agent collision avoidance, then propose a decentralized protocol based on this constraint that leverages gatekeeper to ensure infinite-horizon safety.
III-A R-Boundedness
Our core challenge is to prove network-wide safety using only local information. To solve this, we must first make an agent’s future motion predictable and spatially contained. We achieve this by introducing a geometric constraint called R-Boundedness, which ensures an agent’s entire future trajectory remains within a known, finite volume. This is the critical property that allows agents to formally reason about which other agents they can safely ignore.

### v2

@@ Abstract
Existing decentralized methods for multi-agent motion planning lack formal, infinite-horizon safety guarantees, especially for communication-constrained systems. We present R3R which, to our knowledge, is the first decentralized and asynchronous framework for multi-agent motion planning under range-limited communication constraints with infinite-horizon safety guarantees for systems of nonlinear agents. R3R’s novelty lies in combining our gatekeeper safety framework with a geometric constraint termed R-Boundedness, which together establish a formal link between an agent’s communication radius and its ability to plan safely. We constrain trajectories to lie within a fixed planning radius, determined by a function of the agent’s communication radius. This enables trajectories to be certified as provably safe for all time using only local information. Our algorithm is fully asynchronous, and ensures the forward invariance of these guarantees even in time-varying networks where agents asynchronously join and replan. We evaluate our approach in simulations of up to 128 Dubins vehicles, validating our theoretical safety guarantees in dense, obstacle-rich scenarios. We further show that R3R’s computational complexity scales with local agent density rather than problem size, providing a practical solution for scalable and provably safe multi-agent systems [^1] Code available at github.com/MarshallVielmetti/r3r_decentralized_gatekeeper.
@@ I Introduction
From last mile delivery [1] to warehouse robotics [2] to autonomous drone swarms [3], deploying multi-agent systems safely in the real world is bottlenecked by a lack of formal safety guarantees. This challenge, known as multi-agent motion planning (MAMP), forces a trade-off between guarantees, coordination, and scalability. Centralized planners can provide strong safety guarantees but fail to scale and are prone to single-node failure. Conversely, decentralized methods scale well to large teams but have thus far been unable to provide formal, infinite-horizon safety guarantees under distance-based communication constraints for non-trivial (i.e., beyond single/double integrator) systems.
Multi-agent path finding (MAPF), a discrete state precursor to MAMP, is typically formulated as a graph search problem. Classical MAPF methods, such as Conflict-Based Search and its variants [4, 5, 6, 7], provide strong optimality or bounded suboptimality guarantees, but these approaches are centralized and limited to discrete domains and thus may not produce dynamically feasible solutions. Some MAMP methods have extended these concepts to continuous space using trajectory optimization or sampling-based planning, but these centralized approaches struggle with scalability and real-time performance [8].
@@
Critically, no existing decentralized, asynchronous approach guarantees infinite-horizon safety under distance-based communication constraints. This gap motivates our proposed framework, R3R. To the best of our knowledge, R3R represents the first decentralized and asynchronous framework to provide infinite-horizon safety guarantees for multi-agent motion planning under distance-based communication constraints which generalizes to broad classes of systems under certain assumptions. This is achieved by integrating a novel geometric constraint, which we term R-Boundedness, with gatekeeper [18], our safety framework developed for single agents, to provide recursive feasibility guarantees in multi-agent systems. Our contributions are as follows:
@@
    R3R, a provably safe, decentralized, and asynchronous framework for multi-agent motion planning under communication constraints.
@@
    R-Boundedness, a novel geometric constraint on trajectory generation, which we show links an agent’s ability to communicate with its ability to plan safely.
@@
    We demonstrate the efficacy of our method in simulations of up to 128 agents, showcasing safe and scalable multi-agent coordination.
The paper is organized as follows: Section II introduces preliminaries and the problem statement, section III presents the proposed approach, the methodology is evaluated in simulation in section IV, and conclusions and thoughts for future work are presented in section V.
@@ II Preliminaries and Problem Statement
Notation: $\mathbb{R},\mathbb{R}_{\geq 0},\mathbb{R}_{>0}$ are the sets of reals, non-negative reals, and positive reals. Let $\mathcal{B}_{R}(x)=\{y:\left\|x-y\right\|_{2}\leq R\}$ be the closed ball of radius $R\geq 0$ centered at $x\in\mathbb{R}^{n}$. $\mathcal{A}(t)\subset\mathbb{N}$ denotes the set of all agents at time $t$, where $N=|\mathcal{A}(t)|$ is the total number of agents in the network at that time.
@@ II-A Preliminaries
where $x\in\mathcal{X}\subseteq\mathbb{R}^{n}$ is the state and $u\in\mathcal{U}\subseteq\mathbb{R}^{m}$ is the control input. The function $f:\mathbb{R}_{\geq 0}\times\mathcal{X}\times\mathcal{U}\to\mathbb{R}^{n}$ is piecewise continuous in $t$ and locally Lipschitz in $x$ and $u$. Agents operate in a workspace $\mathcal{W}\subset\mathbb{R}^{d}$, where $d\in\{2,3\}$. Each agent’s state $x$ contains a position component, $x=\begin{bmatrix}p^{\top},\bar{x}^{\top}\end{bmatrix}^{\top}$, where $p\in\mathcal{W}$ is the position and $\bar{x}\in\mathbb{R}^{n-d}$ represents the remaining states. Define the projection $\Pi:\mathcal{X}\to\mathcal{W}$ by $\Pi(x)=p$.
@@
    State $x_{i}$ which evolves through a system satisfying (1).
@@
    Inter-agent avoidance distance of $\delta\in\mathbb{R}_{>0}$, meaning no agent may come within a distance strictly less than $\delta$ of agent $i$.
@@
    Communication radius $R^{\text{comm}}\in\mathbb{R}_{>0}$.
@@
All agents have a common $\delta$ and $R^{\text{comm}}$.
@@
A trajectory [^2] The time-interval $T$ can be a finite horizon, e.g., $T=[t_{0},t_{1}]$ or an infinite horizon, e.g., $T=[t_{0},\infty)$. is the tuple $\mathcal{T}=(T,{\rm x},{\rm u})$, where $T\subset\mathbb{R}_{\geq 0}$ is a time-interval, ${\rm x}:T\to\mathcal{X}$ is the state-trajectory, and ${\rm u}:T\to\mathcal{U}$ is a control-trajectory, satisfying:
$\dot{\rm x}(t)=f(t,{\rm x}(t),{\rm u}(t)),\quad\forall{}t\in T.$ (2)
Definition 3 (Communication Network).
Let $\mathcal{G}(t)=(\mathcal{A}(t),E(t))$ be an undirected, time-varying graph, with nodes representing agents and edges representing communication links, such that:
$E(t)=\{(i,j)\in\mathcal{A}(t):\left\|{\rm p}_{i}(t)-{\rm p}_{j}(t)\right\|_{2}\leq R^{\text{comm}}\}.$ (3)
@@
Communication is instantaneous and range-limited, but not bandwidth-limited. That is, if two agents $i,j$ are neighbors at some time $t$, they can communicate an arbitrary amount of information instantaneously.
Let
$\mathcal{N}_{i}(t)=\{j\in\mathcal{A}(t):(i,j)\in E(t)\}$ (4)
denote the neighbors of agent $i$ at time $t$.
II-B Problem Statement
We now formally state the problem of provably safe multi-agent motion planning under range-limited communication.
Problem 1.
Let $\mathcal{A}(t)\subset\mathbb{N}$ denote a time-varying set of agents satisfying Def.˜1, that communicate over a connectivity graph $\mathcal{G}(t)$ as defined by Def.˜3. Let $\mathcal{W}_{\text{obs}}\subset\mathcal{W}$ be the subset of the domain occupied by static obstacles, and $\mathcal{S}\subset\mathcal{W}\setminus\mathcal{W}_{\text{obs}}$ be the safe set, known to all agents. Agents attempt to navigate to their goal state $g_{i}\in\mathcal{X}$,
$\lim_{t\to\infty}x_{i}(t)=g_{i},\quad\forall i\in\mathcal{A},$ (5)
but must satisfy collision avoidance and safety constraints:
${\rm p}_{i}(t)\in\mathcal{S},$ $\quad\forall t\in\mathbb{R}_{\geq 0},$ (6)
$\left\|{\rm p}_{i}(t)-{\rm p}_{j}(t)\right\|_{2}\geq\delta,$ $\forall j\in\mathcal{A}(t)\setminus\{i\},\forall t\in\mathbb{R}_{\geq 0}.$ (7)
We assume $\mathcal{A}(t_{0})$ is initially empty, and agents are removed from consideration upon reaching their goal state.
To solve ˜1, we propose R3R, a decentralized framework for multi-agent motion planning under communication constraints.
III Proposed Approach and Solution
We construct our proposed approach in three stages. In section˜III-A, we introduce our prior work for single-agent safety gatekeeper [18]. In section˜III-B, we introduce a novel geometric constraint relating agents’ ability to communicate to their ability to plan safely, which we show enables system-wide safety certification using only local information. Finally, in section˜III-C, we integrate these two systems into a provably safe, decentralized algorithm to solve ˜1. An overview of the proposed architecture is shown in Fig.˜1.
@@ III Proposed Approach and Solution
Figure 2: Demonstration of R3R over time for two Dubins agents. The timeline indicates when each agent (Red: Agent 1, Blue: Agent 2) plans or replans. Committed trajectories (solid lines) are certified to be safe for all time, and remain within a planning radius $R^{\text{plan}}$ (filled circles), around where they are constructed. Agents first plan goal-oriented nominal trajectories (dashed lines). A portion of this nominal is combined with a backup trajectory, to form a candidate. If this candidate is deemed safe, it becomes a committed trajectory. At times $t_{0,1}\neq t_{0,2}$ agents $1$ and 22 wish to join the network, then independently construct a valid committed trajectory, which is then simultaneously committed and broadcast as the agent enters the network. (a.) The initial committed trajectories of both agents remain within the planning radius. (b.) Agents continue to track their initial committed trajectories, and enter communication range. (c.) Agent 11 replanned first, so constructs a new candidate trajectory, which it certifies as safe with respect to the current committed trajectory of agent $2$. At time $t_{1_{2}}>t_{1_{1}}$, agent 22 chooses to replan, and must now avoid the recently committed trajectory of agent $1$, which reflects the asynchronous nature of the algorithm; at subsequent replanning steps, agents may not invalidate previously certified decisions of neighbors. (d-f.) The system continues to evolve, and agents continue to asynchronously replan.
III-A gatekeeper Preliminaries
R3R builds upon our prior work for single-agent safety gatekeeper [18], which is executed independently by each agent $i\in\mathcal{A}$ at discrete, asynchronous iterations indexed $k_{i}\in\mathbb{N}$ at times $t_{k_{i}}$, such that $t_{k_{i}}<t_{k_{i}+1}$. We ensure agents always track a committed trajectory that is known to be safe for all future time. At each iteration, agents attempt to construct a new, provably safe, committed trajectory, by properly appending a goal-oriented nominal trajectory with a safe, infinite-horizon backup trajectory. This becomes a candidate trajectory, which must be validated as safe prior to being committed.
Definition 4 (Nominal Trajectory).
Given state ${\rm x}_{i}(t_{k_{i}})\in\mathcal{X}$, at some time $t_{k_{i}}\in\mathbb{R}_{\geq 0}$, the nominal trajectory of agent $i$, denoted
$(T_{k_{i},i}^{\rm nom},{\rm x}_{k_{i},i}^{\rm nom},{\rm u}_{k_{i},i}^{\rm nom}),$ (8)
is defined over the horizon $T_{k_{i},i}^{\rm nom}=[t_{k_{i}},t_{k_{i}}+\Delta t_{H}]$, where $\Delta t_{H}\in\mathbb{R}_{>0}$ is the planning horizon.
A nominal trajectory is not guaranteed to satisfy constraints nor provide any guarantees beyond $\Delta t_{H}$.
@@
A set $\mathcal{C}\subset\mathcal{X}$ is a backup set if $\Pi(\mathcal{C})\subset\mathcal{S}$, and there exists a controller $\pi^{B}:\mathbb{R}\times\mathcal{X}\to\mathcal{U}$ such that $\forall t_{0}\in\mathbb{R}_{\geq 0}$, the closed-loop system $\dot{x}=f(t,x,\pi^{B}(t,x))$ satisfies:
${\rm x}(t_{0})\in\mathcal{C}\implies{\rm x}(t)\in\mathcal{C},\forall t\geq t_{0}.$ (9)
@@
Let $\Delta t_{B}\in\mathbb{R}_{>0}$. For any $t_{0}\in\mathbb{R}_{>0}$, $x_{0}\in\mathcal{X}$, and backup set $\mathcal{C}$, a trajectory $(T^{\rm bak},{\rm x}^{\rm bak},{\rm u}^{\rm bak})$, defined on $T^{\rm bak}=[t_{0},\infty)$ is a backup trajectory if:
${\rm x}^{\rm bak}(t_{0})$ $=x_{0},$ (10)
${\rm p}^{\rm bak}(t)$ $\in\mathcal{S},\;\forall t\in[t_{0},t_{0}+\Delta t_{B}]$ (11)
${\rm x}^{\rm bak}(t_{0}+\Delta t_{B})$ $\in\mathcal{C},$ (12)
and for all $t\geq t_{0}+\Delta t_{B}$, the system evolves under the backup controller
${\rm u}^{\rm bak}(t)=\pi^{B}(t,{\rm x}^{\rm bak}(t)).$ (13)
A backup trajectory takes a system from a state $x_{0}\in\mathcal{X}$ at time $t_{0}$ such that at time $t_{0}+\Delta t_{B}$, the system lies in a backup set. The system then executes the backup controller for all future time, thus remaining in the backup set. If the finite horizon trajectory defined over $t_{0},t_{0}+\Delta t_{B}]$ safely takes the agent into the backup set, then by the properties of the backup set we guarantee the agent will remain in $\mathcal{S}$ over $t_{0},\infty)$.
@@
At time $t_{k_{i}}\in\mathbb{R}_{\geq 0}$, let agent $i\in\mathcal{A}(t_{k_{i}})$ be at state ${\rm x}_{i}(t_{k_{i}})\in\mathcal{X}$. Let the agent’s nominal trajectory be $\mathcal{T}_{k_{i},i}^{\rm nom}=([t_{k_{i}},t_{k_{i}}+\Delta t_{H}],{\rm x}_{{k_{i}},i}^{\rm nom},{\rm u}_{k_{i},i}^{\rm nom})$. A candidate trajectory with switch time $\Delta t_{S}\in[0,\Delta t_{H}]$, denoted $\mathcal{T}_{k_{i},i}^{{\rm can},\Delta t_{S}}=([t_{k_{i}},\infty),{\rm x}_{k_{i},i}^{\rm can},{\rm u}_{k_{i},i}^{\rm can})$, is defined by:
${\rm u}_{k_{i},i}^{\rm can}(\tau)=\begin{cases}{\rm u}_{k_{i},i}^{\rm nom}(\tau)&\text{if }\tau\in[t_{k_{i}},t_{k_{i}}+\Delta t_{S})\\ {\rm u}_{k_{i},i}^{\rm bak}(\tau)&\text{if }\tau\geq t_{k_{i}}+\Delta t_{S}\end{cases},$ (14)
where $([t_{k_{i}}+\Delta t_{S},\infty),{\rm x}_{k_{i},i}^{\rm bak},{\rm u}_{k_{i},i}^{\rm bak})$ is a backup trajectory as defined in Def.˜6, and ${\rm x}_{k_{i},i}^{\rm can}(t)$ is the corresponding state trajectory.
III-B $R$-Boundedness
Our core challenge is to enable system-wide certification of candidate trajectories using local information. To accomplish this, we introduce a geometric constraint called $R$-Boundedness, which restricts an agent to planning within a known, finite volume.